5. The Haber- process has been the primary method for the Industrial production of ammonia (NH 3 ) for over 100 years N 2 (q)+3H 2 (q) 2NH 3 lg) mass of ammonia can be produced by 7.19 * 10 ^ 9 * q of nitrogen gas and 1.19 * 10 ^ 9 / g of hydrogen gas ?
1
Expert's answer
2020-09-29T06:48:07-0400

"N _2 (g)+3H _2 (g) \implies2NH _3 (g)"


moles of Nitrogen here ="\frac{7.19\times10^9}{28}=2.56\times10^8"


moles of Hydrogen here ="\frac{1.19\times10^9}{2}=5.95\times10^8"


here for every mole of nitrogen 3 moles of hydrogen is required and 2 moles of ammonia is formed

clearly from the above mole quantities we can see that hydrogen is the limiting reagent here


"x=\frac{5.95\times10^8}{3}=1.98\times10^8"

moles of ammonia formed ="2x=2\times1.98\times10^8=3.96\times10^8"

mass of ammonia formed ="17\times3.96\times10^8=6.73\times10^9 grams"